OpenStudy (anonymous):

log base4 x- log base 4 (x-1)=1/2 solve the logarithmic function algebraically. round the result to three decimal places.

OpenStudy (cwrw238):

log4 x / (x-1) = 1/2 x / (x-1) = 4^(1/2) = 2 - the rest is algebra are you ok to finish this off?

OpenStudy (anonymous):

i kinda need help with the rest?

OpenStudy (cwrw238):

ok